OpenStudy (anonymous):

they gave me the equation f(x)-f(a)/x-a and f(x)= 5-3x^2. How am I suppose to simplify it.

OpenStudy (lgbasallote):

can you elaborate your question?

OpenStudy (lgbasallote):

how does the equation they gave you look like?

OpenStudy (anonymous):

The equation that they gave me is f(x)= 5-3x^2 and they want me to plug it in for f(x)-f(a)/x-a. Simplify it.

OpenStudy (lgbasallote):

i meant how does f(x) - f(a)/x-a look like?

OpenStudy (lgbasallote):

because im interpreting it as \[\huge f(x) - \frac{f(a)}x - a\]

OpenStudy (anonymous):

its (f(x)-f(a))/(x-a)

OpenStudy (lgbasallote):

so \[\huge \frac{f(x) - f(a)}{x-a}\]

OpenStudy (anonymous):

yes that is the equation...

OpenStudy (lgbasallote):

you have f(x) = 5-3x^2 do you know how to get f(a)?

OpenStudy (anonymous):

I do not know how to get f(a).

OpenStudy (lgbasallote):

hint: substitute x with a

OpenStudy (anonymous):

confuse.

OpenStudy (lgbasallote):

here;s an example \[f(x) = x^2 + 2x + 1\] \[\implies f(a) = a^2 + 2a +1\] i just substituted a with x

OpenStudy (lgbasallote):

got it now?

OpenStudy (anonymous):

so f(a) in this problem will be 5-3a^2?

OpenStudy (lgbasallote):

yes!!

OpenStudy (anonymous):

another question is how do i find x and a

OpenStudy (lgbasallote):

you don't. they stay the same. you just sub in f(x) and f(a)

OpenStudy (anonymous):

so the denominator x-a just stay the same.

OpenStudy (lgbasallote):

yes it does

OpenStudy (anonymous):

I got the answer -3x^2+3a^2. Is that right?

OpenStudy (lgbasallote):

aren't you missing a denominator?

OpenStudy (anonymous):

I do not know how to do the denominator part.

OpenStudy (lgbasallote):

just copy it

OpenStudy (anonymous):

(-3x^2+3a^2)/(x-a)?

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

now factor out -3x^2 + 3a^2

OpenStudy (anonymous):

so it be 3(-x^2+a^2)/(x-a)?

OpenStudy (lgbasallote):

try factoring out -3 instead of 3

OpenStudy (lgbasallote):

what do you get?

OpenStudy (anonymous):

-3(x^2-a^2)/(x-a)

OpenStudy (lgbasallote):

right. now...can you see the difference of two squares tehre?

OpenStudy (anonymous):

(-3)(x+a)(x-a)/(x-a) and the x-a will cancel out right?

OpenStudy (lgbasallote):

yes!!

OpenStudy (anonymous):

will the answer be (-3)(x+a) or do you have to distribute -3 to (x+a)

OpenStudy (lgbasallote):

whatever you please...or if you're in online school...whatever the computer pleases

OpenStudy (anonymous):

If its not online class by keeping it as (-3)(x+a) be correct answer too?

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

okay thanks a lot for the help.

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

I need help how do you find the inverse of f(x)=1+\[\sqrt{1-x^{2}}\]